OpenStudy (anonymous):

1/8^6 rewrite using a negative exponent

OpenStudy (anonymous):

Recall the rule that \(\large\dfrac{1}{a^n} = a^{-n}\). So \(\large \dfrac{1}{8^6} = \ldots\) Do you see what the answer is now? :-)

OpenStudy (anonymous):

8^-6??:)

OpenStudy (anonymous):



OpenStudy (anonymous):

That would be correct! :-)